The capacity of an elevator is 20 children or 15 adults. If 12 children are in the elevator, how many adults can get in?
Answer by edjones(7311)

15/20=3/4 so a kid is comparable to 3/4 of an adult.
so having 12 kids in the elevator is comparible to having 12*3/4=9 adults.
so there is room for 6 adults.
Ed